OpenStudy (anonymous):

(Coordinate Geometry) Please put the equation in standard form. Containing A(5, 3) and perpendicular to a line with slope of -2

OpenStudy (anonymous):

Gradient of line you are looking for is perpendicular to line of slope -2 Thus the gradient is 1/2. Eqn. of line, y= mx +c. Substitute m=1/2 and the point (5,3) 3 = (1/2) 5 +c 3 = 5/2 +c c= 3- 5/2 c= 1/2 Eqn. of line, y= 1/2 x + 1/2